# Amortized Inference of Causal Models via Conditional Fixed-Point Iterations

### v3

@@ 3.3 Inference with Cond-FiP
We provide a summary of inference procedure with Cond-FiP, with details in Section A.3.
Observational Generation. Cond-FiP is capable of generating new data samples: given a random vector noise $\bm{n}\sim\mathbb{P}_{\bm{N}}$, we can estimate the observational sample associated according to an unknown SCM ${\mathcal{S}}(\mathbb{P}_{\bm{N}},{\bm{\mathcal{G}}},\bm{F})\sim{\mathbb{P}}_{{\mathcal{S}}}$ as long as we have access to its empirical representation $(D_{\bm{X}},{\bm{\mathcal{G}}})$. Formally, starting from $\bm{n}_{0}=\bm{n}$, we infer the associated observation by computing for $\ell=1,\dots,d$:
@@
Interventional Generation. Cond-FiP also enables the estimation of interventions given an empirical representation $(D_{\bm{X}},{\bm{\mathcal{G}}})$ of an unkown SCM ${\mathcal{S}}(\mathbb{P}_{\bm{N}},{\bm{\mathcal{G}}},\bm{F})\sim{\mathbb{P}}_{{\mathcal{S}}}$. To achieve this, we start from a noise sample $\bm{n}$, and we generate the associated intervened sample $\widehat{\bm{z}}^{\text{do}}$ by directly modifying the conditional SCM provided by Cond-FiP. More specifically, we modify in place the SCM obtained by Cond-FiP, leading to its interventional version $\mathcal{T}^{\text{do}}(\cdot,D_{\bm{X}},{\bm{\mathcal{G}}})$. Now, generating an intervened sample can be done by applying the loop defined in (5), starting from $\bm{n}$ and using the intervened SCM $\mathcal{T}^{\text{do}}(\cdot,D_{\bm{X}},{\bm{\mathcal{G}}})$ rather than the original one.
@@ 4 Experiments
We begin by describing our experimental setup in Section 4.1, and then present the results of our empirical analysis in Section 4.2, where we benchmark Cond-FiP against state-of-the-art baselines.
@@ 4.1 Setup
Data Generation Process. We use the synthetic data generation procedure proposed by Lorch et al. (2022) to generate SCMs as this framework supports a wide variety of SCMs, making it well-suited for amortized training. It allows sampling of graphs from different schemes and noise variables from diverse distributions. Further, we can also control the complexity of causal mechanisms, choosing between linear (LIN) functions or random fourier features (RFF) for non-linear causal mechanisms. We construct two distribution of SCMs, ${\mathbb{P}}_{\text{IN}}$, and ${\mathbb{P}}_{\text{OUT}}$, which vary based on the choice for sampling causal graphs, noise variables, and causal relationships, see Section B.1 for more details.
@@
Test Datasets. We evaluate the model’s generalization both in-distribution and out-of-distribution by sampling test datasets from ${\mathbb{P}}_{\text{IN}}$ and ${\mathbb{P}}_{\text{OUT}}$, respectively. The test datasets are categorized as follows: LIN IN  and RFF IN  where the SCM are sampled from ${\mathbb{P}}_{\text{IN}}$ with linear and non-linear causal mechanisms respectively. Similarly, we define LIN OUT  and RFF OUT  where the SCMs are sampled from ${\mathbb{P}}_{\text{OUT}}$ instead.
For each category, we vary the total nodes $d\in[10,20,50,100]$ and sample 6 or 9 SCMs per $d$, based on the available schemes for sampling the causal graphs (check Section B.1 for details). This results in a total of $120$ test datasets, supporting a comprehensive evaluation of the methods. For each SCM we generate $n_{\text{test}}=800$ samples, split equally into task context $D_{\bm{X}}$ and queries $D_{\bm{X}^{\prime}}$ for evaluation. An interesting aspect of our test setup is we assess the model’s ability to generalize to larger graphs ($d=50$, $d=100$), despite training only with $d=20$ node graphs.
Model Architecture. For both the dataset encoder and cond-FiP, we set the embedding dimension to $d_{h}=256$ and the hidden dimension of MLP blocks to $512$. Both of our transformer-based models contains 4 attention layers and each attention consists of $8$ attention heads. Please check Section B.2 for further details and Cond-FiP’s memory and compute requirements.
@@
To avoid potential confusion, we clarify that the notion of distribution shift is defined w.r.t Cond-FiP’s training setup. For the baselines, there is no distribution shift as they are trained on the context ($D_{\bm{X}}$) drawn from the specific test distribution. The most important comparison is with the baseline FiP, as Cond-FiP is its amortized counterpart. Further, we do not report detailed comparisons with CausalNF (Javaloy et al., 2023) as its performance was consistently weaker than other baselines, check Appendix J for details.
Evaluation Tasks. We evaluate the methods on the following three tasks. Noise Prediction: given the observations $D_{\bm{X}}$ and the true graph ${\bm{\mathcal{G}}}$, infer the noise variables $\widehat{D_{\bm{N}}}$. Sample Generation: given the noise samples $D_{\bm{N}}$ and the true graph ${\bm{\mathcal{G}}}$, generate the causal variables ^$\widehat{}D_{\bm{X}}$. Interventional Generation: generate intervened samples from noise samples $D_{\bm{N}}$ and the true graph ${\bm{\mathcal{G}}}$.
Metric. Let us denote a predicted & true target as $\widehat{\bm{Y}}\in{\mathbb{R}}^{n_{\text{test}}\times d}$ and $\bm{Y}\in{\mathbb{R}}^{n_{\text{test}}\times d}$. Then RMSE is computed as $\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\sqrt{\frac{1}{d}\|[\bm{Y}]_{i}-[\widehat{\bm{Y}}]_{i}\|_{2}^{2}}$. Note that we scale RMSE by dimension $d$, which allows us to compare results across different graph sizes.
@@
Figure 2: In-Distribution Results. Benchmarking Cond-FiP for various evaluation tasks, with datasets sampled from RFF IN  with $d=20$. The y-axis denotes the RMSE, with mean and standard error over the respective test datasets. Results indicate Cond-FiP can generalize to novel in-distribution instances, with detailed results in Appendix C.
@@
Figure 3: OOD Results. Benchmarking Cond-FiP for various evaluation tasks, with datasets sampled from RFF OUT  with $d=100$ to test for OOD generalization. The y-axis denotes the RMSE, with mean and standard error over the respective test datasets. Results indicate Cond-FiP can generalize to novel OOD instances and larger graphs, with detailed results in Appendix C.
@@
Figure 4: Scarce Data Regime Results. Benchmarking Cond-FiP on the various evaluation tasks (RFF OUT  and $d=100$) as we reduce the test dataset size. The y-axis denotes the RMSE, with mean and standard error over the respective test datasets. Cond-FiP generalizes much better than the baselines in the low-data regime, with detailed results in Appendix E.
@@
Figure 5: OOD Results without True Graph. Benchmarking Cond-FiP for various evaluation tasks, with datasets sampled from RFF OUT  with $d=100$ where the true graph ${\bm{\mathcal{G}}}$ is not present in input context, rather its inferred via AVICI. The y-axis denotes the RMSE, with mean and standard error over the respective test datasets. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph, with detailed results in Appendix F.
@@ 4.2 Results
Generalization to OOD data and larger graphs. In Figure 2, we first present results for in-distribution generalization using test datasets sampled from RFF IN  for graphs with $d=20$ nodes. Cond-FiP performs competitively with baselines trained from scratch on each test instance, hence it successfully generalizes to novel in-distribution instances. Notably, Cond-FiP was never explicitly trained to generate interventional data, and its strong performance on this task further supports that it captures the underlying causal mechanisms.
@@
Next we consider the more challenging case of OOD generalization using test datasets sampled from RFF OUT  and graphs with $d=100$ nodes, while the Cond-FiP was trained only with $d=20$ node graphs. As shown in Figure 3, Cond-FiP continues to perform well, indicating successful generalization to OOD instances and significantly larger graphs! Due to space constraints, we report results for SCMs with non-linear mechanisms—the more challenging setting. Full results for both in-distribution and OOD scenarios are available in Appendix C, where our findings remain consistent.
@@
We also assess Cond-FiP’s sensitivity to distribution shifts by varying the magnitude of distribution shift (details in Appendix D). We consider two cases, where we control the severity in distribution shift by controlling the causal mechanisms or the noise variables. We find that Cond-FiP is more robust to shifts in causal mechanisms, with minimal performance degradation. However, its performance is more sensitive to shifts in noise distributions, deteriorating as the magnitude of shift increases.
@@
Better Generalization in Scare Data Regimes. An advantage of amortized inference methods is their ability to generalize well when context $D_{\bm{X}}$ for test instances is small. As the context size decreases, baselines often suffer significant performance drops as they require training from scratch. In contrast, Cond-FiP is less impacted as its parameters remain unchanged at inference time, and the inductive bias learned during training enables effective generalization even with limited context. In Figure 4, we demonstrate this in the challenging OOD setting (RFF OUT;, $d=100$), where Cond-FiP outperforms the baselines. Please check Appendix E for further details.
@@
Generalization without True Causal Graph. So far, our results assume access to the true causal graph (${\bm{\mathcal{G}}}$) as part of the input context to Cond-FiP. However, Cond-FiP can be extended to operate without this information by first inferring the graph using amortized structure learning methods (Lorch et al., 2022; Ke et al., 2022). We demonstrate this in Figure 5 for the RFF OUT; setting with $d=100$ nodes, using graphs inferred via AVICI (Lorch et al., 2022) for both Cond-FiP and the baselines. The results show that Cond-FiP remains competitive, further supporting its ability to capture underlying causal mechanisms. Please check Appendix F for more details.
@@
Ablation Study. We conduct ablation studies on both the encoder and decoder to better understand how the training data affects generalization performance. We find that Cond-FiP remains competitive even when the encoder is trained on only RFF data, compared to training on a mixture of both. In contrast, decoder performance benefits more noticeably from training on the combined dataset. Please check Section G.1 and G.2 for more details regarding the ablation experiments.
@@
Generalization to novel data simulators. We further evaluate Cond-FiP on test datasets generated using C-Suite (Geffner et al., 2022), a synthetic data simulator distinct from the training simulator. As shown in Figure 6, Cond-FiP generalizes well to these novel instances. Additionally, we consider a modified C-Suite benchmark with Gaussian mixture model noise. Results in Figure 7,  Appendix H show that Cond-FiP also generalizes to instances with more complex noise distributions.
Finally, we show that Cond-FiP can generalize to the real-world instances using the Sachs dataset (Sachs et al., 2005). Although Cond-FiP cannot be trained on real-world datasets since the encoder requires access to true noise variables, it can still be used for inference. We evaluate the quality of generated samples by comparing them to observed data using the Maximum Mean Discrepancy (MMD) metric (Gretton et al., 2012). See Appendix I for more details.
5 Related Works
Amortized Causal Learning.
Amortized inference has gained traction in causality research, particularly for structure learning. Early works by Lorch et al. (2022) and Ke et al. (2022) introduced transformer-based models trained on multiple synthetic datasets using supervised objectives for amortized inference of causal structure. Their approach aligns with recent works on in-context learning of function classes via transformers (Müller et al., 2021; Akyürek et al., 2022; Garg et al., 2022; Von Oswald et al., 2023). Subsequent improvements targeted OOD generalization (Wu et al., 2024), bayesian causal discovery (Dhir et al., 2024), and applications to gene regulatory networks (Ke et al., 2023). Beyond structure learning, amortized methods have been developed for ATE estimation (Nilforoshan et al., 2023; Zhang et al., 2023; Sauter et al., 2025), model selection (Gupta et al., 2023), and partial causal discovery tasks such as learning topological order (Scetbon et al., 2024). However, amortized inference of causal mechanisms in SCMs remains unaddressed, which is the central focus of our work.
Autoregressive Causal Learning. Most causal discovery methods focus first on structure learning (Chickering, 2002; Peters et al., 2014; Zheng et al., 2018), followed by per-node maximum likelihood estimation to recover the causal mechanisms (Blöbaum et al., 2022). In constrast, recent works on causal autoregressive flows (Khemakhem et al., 2021; Geffner et al., 2022; Javaloy et al., 2023) focus on SOTA normalizing flow based generative models to infer causal mechanisms. Further, FiP (Scetbon et al., 2024) modeled SCMs as fixed-point problems over causal (topological) ordering of nodes using transformer-based architectures. These approaches efficiently learn SCMs but require training a separate model per dataset. In this work, we extend FiP to enable amortized inference of causal mechanisms across different SCM instances, removing this limitation.

### v4

@@ 3.3 Inference with Cond-FiP
We provide a summary of inference procedure with Cond-FiP, with details in Section˜A.3.
Observational Generation.
Cond-FiP is capable of generating new data samples: given a random vector noise $\bm{n}\sim\mathbb{P}_{\bm{N}}$, we can estimate the observational sample associated according to an unknown SCM ${\mathcal{S}}(\mathbb{P}_{\bm{N}},{\bm{\mathcal{G}}},\bm{F})\sim{\mathbb{P}}_{{\mathcal{S}}}$ as long as we have access to its empirical representation $(D_{\bm{X}},{\bm{\mathcal{G}}})$. Formally, starting from $\bm{n}_{0}=\bm{n}$, we infer the associated observation by computing for $\ell=1,\dots,d$:
@@
Interventional Generation.
Cond-FiP also enables the estimation of interventions given an empirical representation $(D_{\bm{X}},{\bm{\mathcal{G}}})$ of an unkown SCM ${\mathcal{S}}(\mathbb{P}_{\bm{N}},{\bm{\mathcal{G}}},\bm{F})\sim{\mathbb{P}}_{{\mathcal{S}}}$. To achieve this, we start from a noise sample $\bm{n}$, and we generate the associated intervened sample $\widehat{\bm{z}}^{\text{do}}$ by directly modifying the conditional SCM provided by Cond-FiP. More specifically, we modify in place the SCM obtained by Cond-FiP, leading to its interventional version $\mathcal{T}^{\text{do}}(\cdot,D_{\bm{X}},{\bm{\mathcal{G}}})$. Now, generating an intervened sample can be done by applying the loop defined in (5), starting from $\bm{n}$ and using the intervened SCM $\mathcal{T}^{\text{do}}(\cdot,D_{\bm{X}},{\bm{\mathcal{G}}})$ rather than the original one.
@@ 4.1 Setup
Data Generation Process.
We use the synthetic data generation procedure proposed by Lorch et al. (2022) to generate SCMs as this framework supports a wide variety of SCMs, making it well-suited for amortized training. It allows sampling of graphs from different schemes and noise variables from diverse distributions. Further, we can also control the complexity of causal mechanisms, choosing between linear (LIN) functions or random fourier features (RFF) for non-linear causal mechanisms. We construct two distribution of SCMs, ${\mathbb{P}}_{\text{IN}}$, and ${\mathbb{P}}_{\text{OUT}}$, which vary based on the choice for sampling causal graphs, noise variables, and causal relationships, see Section˜B.1 for more details.
@@
Test Datasets. We evaluate the model’s generalization both in-distribution and out-of-distribution by sampling test datasets from ${\mathbb{P}}_{\text{IN}}$ and ${\mathbb{P}}_{\text{OUT}}$, respectively. The test datasets are categorized as follows: LIN IN  and RFF IN  where the SCM are sampled from ${\mathbb{P}}_{\text{IN}}$ with linear and non-linear causal mechanisms respectively. Similarly, we define LIN OUT  and RFF OUT  where the SCMs are sampled from ${\mathbb{P}}_{\text{OUT}}$ instead. For each category, we vary the total nodes $d\in[10,20,50,100]$ and sample 6 or 9 SCMs per $d$, based on the available schemes for sampling the causal graphs (check Section˜B.1 for details). This results in a total of $120$ test datasets, supporting a comprehensive evaluation of the methods. For each SCM we generate $n_{\text{test}}=800$ samples, split equally into task context $D_{\bm{X}}$ and queries $D_{\bm{X}^{\prime}}$ for evaluation. An interesting aspect of our test setup is we assess the model’s ability to generalize to larger graphs ($d=50$, $d=100$), despite training only with $d=20$ node graphs.
Model Architecture. For both the dataset encoder and cond-FiP, we set the embedding dimension to $d_{h}=256$ and the hidden dimension of MLP blocks to $512$. Both of our transformer-based models contains 4 attention layers and each attention consists of $8$ attention heads. Please check Section˜B.3 for further details and Cond-FiP’s memory and compute requirements.
@@
To avoid potential confusion, we clarify that the notion of distribution shift is defined w.r.t Cond-FiP’s training setup. For the baselines, there is no distribution shift as they are trained on the context ($D_{\bm{X}}$) drawn from the specific test distribution. The most important comparison is with the baseline FiP, as Cond-FiP is its amortized counterpart. Further, we do not report detailed comparisons with CausalNF (Javaloy et al., 2023) as its performance was consistently weaker than other baselines, check Appendix˜H for details.
Evaluation Tasks.
We evaluate the methods on the following three tasks. Noise Prediction: given the observations $D_{\bm{X}}$ and the true graph $\bm{\mathcal{G}}$}, infer the noise variables $\widehat{D_{\bm{N}}}$. Sample Generation: given the noise samples $D_{\bm{N}}$ and the true graph $\bm{\mathcal{G}}$}, generate the causal variables $\widehat{}D_{\bm{X}}$. Interventional Generation: generate intervened samples from noise samples $D_{\bm{N}}$ and the true graph $\bm{\mathcal{G}}$}.
Metric.
Let us denote a predicted & true target as $\widehat{\bm{Y}}\in{\mathbb{R}}^{n_{\text{test}}\times d}$ and $\bm{Y}\in{\mathbb{R}}^{n_{\text{test}}\times d}$. Then RMSE is computed as $\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\sqrt{\frac{1}{d}\|[\bm{Y}]_{i}-[\widehat{\bm{Y}}]_{i}\|_{2}^{2}}$. Note that we scale RMSE by dimension $d$, which allows us to compare results across different graph sizes.
@@
Figure 2: In-Distribution Results. Benchmarking Cond-FiP for various evaluation tasks, with datasets sampled from RFF IN  with $d=20$. The y-axis denotes the RMSE, with mean and standard error over the respective test datasets. Results indicate Cond-FiP can generalize to novel in-distribution instances, with detailed results in Appendix˜C.
@@
Figure 3: OOD Results. Benchmarking Cond-FiP for various evaluation tasks, with datasets sampled from RFF OUT  with $d=100$ to test for OOD generalization. The y-axis denotes the RMSE, with mean and standard error over the respective test datasets. Results indicate Cond-FiP can generalize to novel OOD instances and larger graphs, with detailed results in Appendix˜C.
@@
Figure 4: Scarce Data Regime Results. Benchmarking Cond-FiP on the various evaluation tasks (RFF OUT  and $d=100$) as we reduce the test dataset size. The y-axis denotes the RMSE, with mean and standard error over the respective test datasets. Cond-FiP generalizes much better than the baselines in the low-data regime, with detailed results in Appendix˜E.
@@
Figure 5: OOD Results without True Graph. Benchmarking Cond-FiP for various evaluation tasks, with datasets sampled from RFF OUT  with $d=100$ where the true graph $\bm{\mathcal{G}}$} is not present in input context, rather its inferred via AVICI. The y-axis denotes the RMSE, with mean and standard error over the respective test datasets. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph, with detailed results in Appendix˜F.
@@
(a) Noise Prediction
(b) Sample Generation
(c) Interventional Generation
Figure 7: CSuite GMM Results. Benchmarking Cond-FiP on the Large Backdoor and Weak Arrow datasets from the CSuite benchmark, where the noise distribution is modified to be a multi-modal gaussian mixture model. The y-axis denotes the RMSE, with mean and standard error across the 12 test scenarios. Results indicate that Cond-FiP can generalize to instances with more complex noise distributions like GMMs.
@@ 4.2 Results
In Figure˜2, we first present results for in-distribution generalization using test datasets sampled from RFF IN  for graphs with $d=20$ nodes. Cond-FiP performs competitively with baselines trained from scratch on each test instance, hence it successfully generalizes to novel in-distribution instances. Notably, Cond-FiP was never explicitly trained to generate interventional data, and its strong performance on this task further supports that it captures the underlying causal mechanisms.
@@
Next we consider the more challenging case of OOD generalization using test datasets sampled from RFF OUT  and graphs with $d=100$ nodes, while the Cond-FiP was trained only with $d=20$ node graphs. As shown in Figure˜3, Cond-FiP continues to perform well, indicating successful generalization to OOD instances and significantly larger graphs! Due to space constraints, we report results for SCMs with non-linear mechanisms—the more challenging setting. Full results for both in-distribution and OOD scenarios are available in Appendix˜C, where our findings remain consistent.
@@
We also assess Cond-FiP’s sensitivity to distribution shifts by varying the magnitude of distribution shift (details in Appendix˜D). We consider two cases, where we control the severity in distribution shift by controlling the causal mechanisms or the noise variables. We find that Cond-FiP is more robust to shifts in causal mechanisms, with minimal performance degradation. However, its performance is more sensitive to shifts in noise distributions, deteriorating as the magnitude of shift increases.
Better Generalization in Scarce Data Regimes.
@@
An advantage of amortized inference methods is their ability to generalize well when context $D_{\bm{X}}$ for test instances is small. As the context size decreases, baselines often suffer significant performance drops as they require training from scratch. In contrast, Cond-FiP is less impacted as its parameters remain unchanged at inference time, and the inductive bias learned during training enables effective generalization even with limited context. In Figure˜4, we demonstrate this in the challenging OOD setting (RFF OUT , $d=100$), where Cond-FiP outperforms the baselines .Please check Appendix˜E for further details.
Generalization without True Causal Graph.
@@
So far, our results assume access to the true causal graph ($\bm{\mathcal{G}}$}) as part of the input context to Cond-FiP. However, Cond-FiP can be extended to operate without this information by first inferring the graph using amortized structure learning methods (Lorch et al., 2022; Ke et al., 2022). We demonstrate this in Figure˜5 for the RFF OUT; setting with $d=100$ nodes, using graphs inferred via AVICI (Lorch et al., 2022) for both Cond-FiP and the baselines. The results show that Cond-FiP remains competitive, supporting its ability to capture underlying causal mechanisms (details in  Appendix˜F).
Further, to assess whether Cond-FiP genuinely learns causal mechanisms tied to the input graph structure, we perform a systematic sensitivity analysis by perturbing the causal graph. Starting from the true graph, we randomly remove a proportion ($p$) of the true edges, such that on average $p\times$ (total edges) are missing. As shown in Table 21 (Appendix F.1), Cond-FiP remains robust under moderate graph perturbations and performs competitively with FiP, which is retrained from scratch for each setting. These results further support that Cond-FiP learns transferable causal mechanisms and can adapt at test time by inferring functions consistent with the available (and potentially imperfect) graph and observational context.
Ablation Study.
Since the method relies on large-scale pretraining over synthetic SCMs, we analyze how Cond-FiP’s performance scales with the number of pretraining SCMs. We conduct experiments at smaller scales, with a total of $1e5,4e5,1e6$ SCMs, as opposed to using $4e6$ SCMs in our main results above. Our results in Table 22 (Section˜G.1) show that Cond-FiP benefits consistently from additional pretraining data, though the returns gradually diminish as scale increases. The most pronounced gains appear when increasing the pretraining size from $1e5$ to $4e5$, while improvements become more incremental beyond $1e6$.
@@
Further, we conduct ablation studies on both the encoder (Section˜G.2) and decoder (Section˜G.3) to better understand how the training data affects generalization performance. We find that Cond-FiP remains competitive even when the encoder is trained on only RFF data, compared to training on a mixture of both. In contrast, decoder performance benefits more noticeably from training on the combined dataset.
Generalization to novel data simulators.
@@
We further evaluate Cond-FiP on test datasets generated using C-Suite (Geffner et al., 2022), a synthetic data simulator distinct from the training simulator. As shown in Figure˜6, Cond-FiP generalizes well to these novel instances. Additionally, to conduct more OOD evaluations, we modify the noise distribution of the Large Backdoor and Weak Arrow datasets from the Csuite benchmark such that the noise variables are sampled from a gaussian mixture model (GMM) (details in Section˜B.2). Results in Figure˜7 demonstrate that Cond-FiP can generalize to more complex noise distributions as well. Importantly, while baselines were trained from scratch for each specific gaussian mixture noise distribution, Cond-FiP was pretrained only on gaussian noise and generalizes effectively to settings with GMM noise distribution.
DoWhy $0.015$ $0.014$ $0.005$
DECI $0.014$ $0.005$ $0.005$
FiP $0.015$ $0.005$ $0.005$
Cond-FiP $0.013$ $0.005$ $0.005$
Table 1: Results for Flow Cytometry (Sachs) dataset. We benchmark Cond-FiP against the baselines for the task of generating observational data on the real world Sachs benchmark. Each cell reports the MMD, and we also report the reconstruction error for all of the methods. Results indicate that Cond-FiP matches the performance of baselines trained from scratch.
DoWhy $0.020$ $0.014$ $0.005$
DECI $0.016$ $0.005$ $0.005$
FiP $0.017$ $0.005$ $0.005$
Cond-FiP $0.019$ $0.005$ $0.005$
Table 2: Results for Ecoli dataset. We benchmark Cond-FiP against the baselines for the task of generating observational data on the real world Ecoli benchmark from the bnlearn repository. Each cell reports the MMD, and we also report the reconstruction error for all of the methods. Results indicate that Cond-FiP matches the performance of baselines trained from scratch.
Experiments on real-world benchmarks.
Finally, we show that Cond-FiP can generalize to the real-world instances using the flow cytometry dataset (Sachs et al., 2005) and ecoli dataset (Scutari, 2010). Although Cond-FiP cannot be trained on real-world datasets since the encoder requires access to true noise variables, it can still be used for inference. Both datasets contains $n\simeq 800$ observational samples expressed in a $d=11$ dimensional space for the flow cytometry dataset and $d=46$ dimensional space for the ecoli dataset, and the corresponding reference (true) causal graph. We split this into context $D^{\text{context}}_{\bm{X}}\in{\mathbb{R}}^{n_{\text{context}}\times d}$ and queries $D^{\text{query}}_{\bm{X}}\in{\mathbb{R}}^{n_{\text{query}}\times d}$, each of size $n_{\text{context}}=n_{\text{query}}=400$. Note that the context dataset is to used to train the baselines and obtain dataset embedding for Cond-FiP, while the query dataset is used for evaluation of all the methods.
Since we don’t have access to the true causal mechanisms, we cannot compute RMSE for noise prediction or sample generation like we did in our experiments with synthetic benchmarks. Instead for each method, we obtain the noise predictions $\widehat{D^{\text{context}}_{\bm{N}}}$ on the context, and use it to fit a gaussian distribution for each component (node). Then we use the learned gaussian distribution to sample new noise variables, $\widehat{D^{\text{query}}_{\bm{N}}}$, which are mapped to the observations as per the causal mechanisms learned by each method, $\widehat{D^{\text{query}}_{\bm{X}}}$. Finally, we compute the maximum mean discrepancy (MMD) distance between $\widehat{D^{\text{query}}_{\bm{X}}}$ and $D^{\text{query}}_{\bm{X}}$ as metric to determine whether the method has captured the true causal mechanisms. For consistency, we also evaluate the reconstruction performances by directly using the inferred noise from context $\widehat{D^{\text{context}}_{\bm{N}}}$ from the models, and then compute MMD between their reconstructed data ($\widehat{D^{\text{context}}_{\bm{X}}}$) and the query data ($D^{\text{query}}_{\bm{X}}$).
Table 1,  2, presents our results for the flow cytometry and the ecoli dataset, where for reference we also report the MMD distance between samples from the context and query split, which should serve as the gold standard since both the datasets are sampled from the same distribution. We find that Cond-FiP is competitive with the baselines that were trained from scratch. Except DoWhy, the MMD distance with reconstructed samples from the methods are close to oracle performance.
Note that Cond-FiP (as well as the other baselines) only supports hard interventions while the interventional data available for Sachs are soft interventions. Hence, we are unable to provide a comprehensive evaluation of Cond-FiP (and the baselines) for interventional predictions on Sachs.
Computational Efficiency.
For instance, Cond-FiP can infer causal mechanisms for a novel task in under one minute, while FiP takes on average $30$ minutes per task. For a concrete comparison, it took us $30$ hours to train Cond-FiP but we can solve each inference task in max $1$ minute. Therefore, to compute our main results ( Appendix˜C we evaluated $360$ different tasks, implying a total cost of $30+360/60=36$ hours with Cond-FiP. In contrast, evaluating FiP on the same $360$ tasks would require retraining from scratch each time, taking approximately $180$ hours!
Thus, while Cond-FiP has a higher one-time training cost, it offers a $5\times$ speedup over FiP in total runtime when evaluating across multiple tasks.